OpenStudy (anonymous):

joshni threw a ball and its height is represented by the function h(t) = -16t2 + 30t + 5. Explain how she can determine the maximum height the ball reached. Use complete sentences to receive full credit.

OpenStudy (anonymous):

do you know how to find vertex?

OpenStudy (anonymous):

since the height is a parabola, its maximum point is the vertex of the parabola.

OpenStudy (anonymous):

a downward parabola*

OpenStudy (anonymous):

parabola : ax^2 + bx + c vertex formula = -b/ (2a) vertex = -30/(2*-16) vertex = ? lemme know if you need more hints.